Shift Summary for DAY, Dec. 25, 1999

Angela Bellavance (bellavan@rice.edu)
Sat, 25 Dec 1999 15:36:53 -0600 (CST)

Shift Summary for day shift, 25/Dec/99

On Shift: P. Shanahan, A. Bellavance

Hours of beam delivered at all intensities: 4.3
Hours of beam delivered at nominal intensity: 3.9
Hours of E799 data (special runs not included): 3.4
Hours of special runs (laser scans, muon runs...): 0.8

Accounting of lost time: TeV quenches and other TeV problems.
Our messing with the kumquats / ktrig window.

========================================================================
Runs ended this shift:
Run 15286 - 12.9M events - BEAM_799
Run 15287 - LASERSCAN
Run 15288 - 2.6M events - BEAM_799
Run 15289 - ongoing

Tapes brought to FCC this shift: UPA253-258
UPB253-258
UPC253-258
UPS085-86
UPF064
UPH046
UPM049
UPPE10

========================================================================
Accesses during shift: none
========================================================================
Overview of shift:
Smooth data taking through early morning. Beam disappears around 9 and
doesn't return until 13:00. We crash ourselves in the afternoon and miss
0.9 hours of beam.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15286 07:00 Take over, run ongoing.
07:46 Repeated request for tape mount of UPLA12 in
ktev3-tps2d4. Arun e-mailed about possibly faulty
drive. We clean the drive again and try to
re-mount the tape again. It looks like it's being
written to...maybe a software fault?
212 08:41 Reload Etotal and HCC to try to fix a small
inefficiency.
221 08:52 TeV abort and we enter a SY conservation period.
237 09:13 HCC ineff. alarm. Goes away next spill on its
own. No beam yet.
09:43 Message from web page: left bends trip.
09:50 TeV quench at B4.
09:56 Ch. 13:"Est. 15 minutes."
10:25 Another TeV quench at B4.
10:39 Ch. 13:"Est. 20 minutes."
11:18 Ch. 13:"Estimate 45 minutes."
11:28 Stop run to do a laser scan. Run refuses to
finish. Do fresh-start.
11:45 UPLA wants to be mounted again. Arun paged. He's
coming in for a look.
15287 12:01 Start Laser scan.
15288 1 13:00 Start run. Beam returns on spill 2. Nominal
intensity by spill 5.
11-13 13:13 TeV aborts.
19 MCR warns us they are tuning. Lots of scaler
alarms.
44 MCR done tuning. # scaler alarms much less now.
14:06 Tape mount request for UPK034.
58 14:18 Looking at histograms discovers a problem with the
kumquats for chamber 3. Try reloading kumquats.
60 Reload didn't work. Try a reset of the crate.
That doesn't work, send for Tony.
14:45 Lab Director and Mrs. Witherell stop by KTeV with
a delivery of home-made chocolate chip cookies!
They are heartfully appreciated.
15:05 Run Control appears hopelessly hung. Do a
drt-fresh-start and load new tapes.
15289 1 15:11 Run started successfully. And the original
problem is still there! Tony says don't worry
about it.

========================================================================
Follow up Notes on Previous Shift

Don't worry about the "heater" temp. Just be sure the "CsI-Bot" and
CsI-Top" readings are about 15 and 19 degrees, respectively.

========================================================================
Items to Follow up on Next Shift

-kumquat for chamber 3 has a hot channel. Wait for expert to return from
vacation.
-inefficient HCC bit 1767. Bad connection in ET that will be worked on
next long down time.
-latches are out of the readout so some of the monitoring histograms look
funny.

_______
Angela Bellavance work: (630) 840-2801
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke